### v1

@@ Abstract
Context. FRB 20121102A was the first fast radio burst to be observed to repeat. Since then, thousands of bursts have been detected by multiple radio telescopes around the world. Previous work has shown an indication of a cyclic activity level with a periodicity around 160 days. Knowing when the source repeats is essential for planning multi-wavelength monitoring to constrain their emission extend and progenitor source.
@@
Methods. We use the Lomb-Scargle periodogram on a sample of 272 observing epochs where correspond to detections and to non-detections. Our dataset is composed of the 7 epochs of our monitoring plus publicly available data. We investigate two methods, i) binary model, describing the observing epochs with if there are detections and with 0 for non-detections. ii) normalised rates model: which considers the inferred detections rates.
Results. We report no detections in 12.5-hour observations down to a fluence of . The best period found for the cyclic activity window is days for the binary model and days for the normalised rates model. The activity phase is shown to be . The normalised rates shows a clear Gaussian-like behaviour for the activity level, where the number of detections peak at the centre of the activity window.
@@
. Start date (UTC) Duration () Sample time () Number of channels 2022-09-18 10:08:48 2022-09-18 23:58:13 2022-11-19 23:10:17 2023-02-25 18:53:49 2023-06-29 14:54:23 2023-08-19 03:10:42 2023-11-17 01:41:03
@@ 2 Follow-up
The observations were carried out with the central beam of the 7-beam receiver (a description of this can be found in Cruces et al. 2021a) in combination with Effelsberg’s direct digitalization system (EDD) with a bandwidth at central frequency of with several time and frequency resolutions, going from and frequency channels, to and frequency channels in the largest dataset.
@@ 4.1 Non-detections
No bursts were detected in of data of FRB 20121102A at L-band frequencies taken with Effelsberg down to a S/N of 7. This means, no detections down to a fluence111In this work we compute fluence based on the band-averaged peak flux density. of considering the average of the pulse width222We refer to the band-averaged pulse width as pulse width. from the bursts reported in Cruces et al. (2021b). This fluence corresponds to an isotropic energy of .
@@ 4.2.1 L-band dataset
The dataset is made from the observations reported in Spitler et al. (2014); Scholz et al. (2016); Spitler et al. (2016); Scholz et al. (2017); Houben et al. (2019); Oostrum et al. (2020); Rajwade et al. (2020); Caleb et al. (2020); Li et al. (2021); Cruces et al. (2021b); Hewitt et al. (2022); Jahns et al. (2023); Feng et al. (2023) and the observations reported in this work. We made sure to only include each observation once, as some were reported more than once by different authors. In case of only reported modified Julian date (MJD) of detected bursts we used the MJDs of the bursts and the wait times between them to determine approximate start of observation times. We included in the dataset a single burst reported in Feng et al. (2023) because there is a lack of detections between and . The full dataset consists of epochs in a total time span of days with non-detections and detections. We use the Lomb-Scargle periodogram (VanderPlas, 2018) technique to find the periodicity of the source and find a periodicity of days which is in very good agreement with the days found in Cruces et al. (2021b).
@@
Along with the binary detection model, we used a more sophisticated model in which we instead use the rates of detections for each observation in our dataset. To address the difference in sensitivity across the telescopes, we translate the rates to what a 100-m telescope would infer. We refer to this rate as normalised rate, and we calculate it as:
@@
where the S/N is taken to be , is the number of polarisations taken to be 2, is the pulse width considered to be , and is the observation bandwidth.
@@ 5 Discussion
The results obtained for the periodicity of FRB 20121102A are consistent and further improved than the previously reported in Cruces et al. (2021b) because our dataset contains more epochs and an extended time span. We tried the method of normalising the rates to test whether it improves the estimation of the periodicity of FRB 20121102A and find a period of days. In comparison with the days obtained from the binary model, both peaks have a significance greater than but the normalised rate model has lower significance of the binary model. We attribute the difference to the fact that in the binary model the power of the signal is split between detections (1) and non-detections (0), while in the normalised rate model the power is distributed in a wider bin range. This introduces as well more noise to the timeseries. A more accurate model to normalise the rates would consider different values in eq. 2 for each telescope in the dataset. To test the importance of the parameter in our results, aside from the obtained in Cruces et al. (2021b) we try using reported in Li et al. (2021) and obtain a consistent periodicity. Therefore because we used Effelsberg as our reference in the normalisation we consider that using for all telescopes is optimal. Because of the higher significance of the peak in the binary model, we take this as the new best periodicity for FRB 20121102A. To centre the activity window around a phase of , we use the MJD as a reference epoch for phase and we get an activity window of . This means, that roughly out of the days duration of one cycle, the source is active within a days window at L-band. This is broader than the % for FRB 20180916B (CHIME/FRB Collaboration et al., 2020).
@@
Regarding the dependence of the activity window with the observing frequency, we obtain for the L and S-band the same period. This means that the phenomena leading to the active phase of the source is intrinsic. However, as is seen in the folding of the S-band dataset shown in Fig. 4, the full phase range is not mapped. In particular we lack observations in a considerable part of the phase, towards the start of the activity. We can see that the S-band normalised rates shows as well some indication of a Gaussian-like profile for the activity window but is not possible to infer its duration as it is the case for the L-band dataset. Folding the S-band dataset with the periodicity obtained at L-band we find that the activity window for S-band ends before the L-band activity window, having a phase difference of 0.12. This means that the S-band activity phase finishes 19 days before the L-band active phase. Since we do not have the part of the phase completely mapped in S-band we cannot conclude for certain what happens in this part.
Regarding the C-band dataset, we do not have enough observations and in particular detections for the Lomb-Scargle periodogram to output a consistent periodicity. More observations will help better understand the shift in phase between frequency bands and the behaviour of the activity phase (Espinoza et al. in prep.).
@@ 6 Conclusions
We reported 7 observing epochs of FRB 20121102A monitoring and find no detections. We combined our observations with publicly available data of the source and divided it per frequency band and found a new best periodicity for the cyclic activity window to be days at L-band. This updated periodicity is more precise and with higher significance than the previously reported in Cruces et al. (2021a), mainly because of the larger and extended dataset. We also report the same periodicity using a normalised rates model that is more precise but has less significance than the one obtained by the binary model. We found a consistent periodicity at S-band and report that this active phase seems to end 19 days before the L-band active phase. No conclusions can be drawn from the C-band observations given the lack the observing epochs and in particular of detections. We strongly encourage the community to report both detections and non-detections of repeating FRBs and the full observation details. Sky and instrumentation statistics such as exposure time, observation duration, and telescope properties are fundamental to further constrain the nature of FRBs.

### v2

@@ Abstract
Context. FRB 20121102A was the first fast radio burst to be observed to repeat. Since then, thousands of bursts have been detected by multiple radio telescopes around the world. Previous work has shown an indication of a cyclic activity level with a periodicity around 160 days. Knowing when the source repeats is essential for planning multi-wavelength monitoring to constrain their emission extent and progenitor source.
@@
Methods. We use the Lomb-Scargle periodogram on a sample of 284 observing epochs where correspond to detections and to non-detections. Our dataset is composed of the 7 epochs of our monitoring plus publicly available data. We investigate two methods, i) binary model, describing the observing epochs with if there are detections and with 0 for non-detections. ii) normalised rates model: which considers the inferred detections rates.
Results. We report no detections in 12.5-hour observations down to a fluence of . The best period found for the cyclic activity window is days for the binary model and days for the normalised rates model. The activity phase is shown to be . The normalised rates show a clear Gaussian-like behaviour for the activity level, where the number of detections peak at the centre of the activity window.
@@
Start date (UTC) | Duration () | Sample time () | Number of channels |
2022-09-18 10:08:48 | |||
2022-09-18 23:58:13 | |||
2022-11-19 23:10:17 | |||
2023-02-25 18:53:49 | |||
2023-06-29 14:54:23 | |||
2023-08-19 03:10:42 | |||
2023-11-17 01:41:03 |
@@ 2 Follow-up
The observations were carried out with the central beam of the 7-beam receiver (a description of this can be found in Cruces et al. 2021a) in combination with Effelsberg’s direct digitalisation system (EDD) with a bandwidth at central frequency of with several time and frequency resolutions, going from and frequency channels, to and frequency channels in the largest dataset.
@@ 4.1 Non-detections
No bursts were detected in of data of FRB 20121102A at L-band frequencies taken with Effelsberg down to a S/N of 7. This means, no detections down to a fluence222In this work we compute fluence based on the band-averaged peak flux density. of considering the average of the pulse width333We refer to the band-averaged pulse width as pulse width. from the bursts reported in Cruces et al. (2021b). This fluence corresponds to an isotropic energy of .
@@ 4.2.1 L-band dataset
The dataset is made from the observations reported in Spitler et al. (2014); Scholz et al. (2016); Spitler et al. (2016); Scholz et al. (2017); Hardy et al. (2017); Marcote et al. (2017); Houben et al. (2019); Oostrum et al. (2020); Rajwade et al. (2020); Caleb et al. (2020); Li et al. (2021); Cruces et al. (2021b); Hewitt et al. (2022); Jahns et al. (2023); Feng et al. (2023); Zhang et al. (2024) and the observations reported in this work. The dataset contains observations made with Effelsberg, the Arecibo Observatory with the Arecibo L-Band Feed Array (ALFA) and L-wide receivers, the Westerbork Synthesis Radio Telescope (WSRT), the Five-hundred-meter Aperture Spherical radio Telescope (FAST), the Green Bank Telescope (GBT), the Lovell Telescope (LT), the Meerkat Telescope, the Very Large Array (VLA) and the Deep Space Network telescopes 43 and 63 (DSS-43 and DSS-63). We made sure to only include each observation once, as some were reported more than once by different authors. The timestamps used correspond to the modified Julian date (MJD) marking the start of each observation. In case of only reported MJDs of detected bursts we used the MJDs of the bursts and the wait times between them to determine approximate start of observation times. We included in the dataset a single burst reported in Feng et al. (2023) because there is a lack of detections between and . The full dataset consists of epochs in a total time span of days with non-detections and detections. We use the Lomb-Scargle periodogram (VanderPlas, 2018) technique to find the periodicity of the source and find a periodicity of days which is in very good agreement with the days found in Cruces et al. (2021b).
@@
Along with the binary detection model, we used a more sophisticated model in which we instead use the rates of detections, calculated as the number of detected bursts within the observation length for each observation in our dataset. To address the difference in sensitivity across the telescopes, we translate the rates to what a 100-m telescope would infer. We refer to this rate as normalised rate, and we calculate it as:
@@
where the S/N is taken to be , is the number of polarisations taken to be 2, is the pulse width considered to be , and is the observation bandwidth. One bias in this model is that, although the detections are weighted relative to each other by the event rate, all non-detections are weighted equally, regardless of their duration. To mitigate this, we have split the non-detections into 1-hour blocks (each representing 0 events/hour). We removed the blocks of non-detections that were left with a duration of less than 1-hour after the split. We consider the bias in this approach less significant than treating each non-detection observation as an independent data point, regardless of its duration.
@@ 5 Discussion
The results obtained for the periodicity of FRB 20121102A are consistent and further improve the previously reported in Cruces et al. (2021b) because our dataset contains more epochs and an extended time span. We tried the method of normalising the rates to test whether it improves the estimation of the periodicity of FRB 20121102A and find a period of days. The advantage of the normalised rate model is that it weighs each timestamp by the level of activity within it. In comparison with the days obtained from the binary model, both peaks have a significance greater than but the normalised rate model has a lower significance than the binary model. We attribute the difference to the fact that in the binary model the power of the signal is split between detections (1) and non-detections (0), while in the normalised rate model the power is distributed in a wider bin range. This introduces as well more noise to the timeseries. An in-depth analysis of the different methods to estimate the periodicity is work in progress.
Additionally, a more accurate model for the normalised rates would consider different values in eq. 2 for each telescope in the dataset. To test the importance of the parameter in our results, aside from the obtained in Cruces et al. (2021b) we try using reported in Li et al. (2021) and obtain a consistent periodicity. Therefore because we used Effelsberg as our reference in the normalisation we consider that using for all telescopes is optimal. Because of the higher significance of the peak in the binary model, we take this as the new best periodicity for FRB 20121102A. To centre the activity window around a phase of , we use the MJD as a reference epoch for phase and we get an activity window of . This means, that roughly out of the days duration of one cycle, the source is active within a days window at L-band. This is broader than the % for FRB 20180916B (CHIME/FRB Collaboration et al., 2020).
@@
Regarding the dependence of the activity window with the observing frequency, we obtain for the L and S-band a consistent period. For the normalised rate model the 163.1 days value obtained for the periodicity is consistent with both the S-band binary model and with the L-band value within . This means that the phenomena leading to the active phase of the source is intrinsic. However, as is seen in the folding of the S-band dataset shown in Fig. 4, the full phase range is not mapped. In particular we lack observations in a considerable part of the phase, towards the start of the activity. We do not see a clear gaussian-like trend as it is appreciated in the L-band dataset, probably because of the limited dataset. More observations in this band will permit a quantitative comparison across frequencies.
Folding the S-band dataset with the periodicity obtained at L-band we find that the activity window for S-band ends before the L-band activity window, having a phase difference of 0.12. This means that the S-band activity phase finishes 19 days before the L-band active phase. Since we do not have the part of the phase completely mapped in S-band we cannot conclude for certain what happens in this part.
Regarding the C-band dataset, we do not have enough observations and in particular detections for the Lomb-Scargle periodogram to output a consistent periodicity. More observations will help better understand the shift in phase between frequency bands and the behaviour of the activity phase (Espinoza-Dupouy et al. in prep.).
@@ 6 Conclusions
We reported 7 observing epochs of FRB 20121102A monitoring and find no detections. We combined our observations with publicly available data of the source and divided it per frequency band and found a new best periodicity for the cyclic activity window to be days at L-band. This updated periodicity is more precise and with higher significance than the previously reported in Cruces et al. (2021b), mainly because of the larger and extended dataset. We also report the same periodicity using a normalised rates model that is more precise but has less significance than the one obtained by the binary model. We found a consistent periodicity at S-band and report that this active phase seems to end 19 days before the L-band active phase. No conclusions can be drawn from the C-band observations given the lack the observing epochs and in particular of detections. We strongly encourage the community to report both detections and non-detections of repeating FRBs and the full observation details. Sky and instrumentation statistics such as exposure time, observation duration, and telescope properties are fundamental to further constrain the nature of FRBs.
@@ Acknowledgements.
- Zhang et al. (2024) Zhang, Y.-K., Li, D., Feng, Y., et al. 2024, Science Bulletin, 69, 1020
Telescope | (Jy ms) | Frequency Band |
Effelsberg | 0.15 | L-band |
Arecibo | 0.03 / 0.02 | L-band (ALFA / L-Wide) |
GBT | 0.05 | S-band |
FAST | 0.012 | L-band |
WSRT | 0.26 | L-band |
Meerkat | 0.06 | L-band |
DSS-43 | 0.23 | S-band |
DSS-63 | 0.5 | S-band |
VLA | 0.19 / 0.06 | L-band / S-band |
Lovell | 0.23 | L-band |